### v2

# 4DHands: Reconstructing Interactive Hands in 4D with Transformers

Abstract
In this paper, we introduce 4DHands, a robust approach to recovering interactive hand meshes and their relative movement from monocular inputs. Our approach addresses two major limitations of previous methods: lacking a unified solution for handling various hand image inputs and neglecting the positional relationship of two hands within images. To overcome these challenges, we develop a transformer-based architecture with novel tokenization and feature fusion strategies. Specifically, we propose a Relation-aware Two-Hand Tokenization (RAT) method to embed positional relation information into the hand tokens. In this way, our network can handle both single-hand and two-hand inputs and explicitly leverage relative hand positions, facilitating the reconstruction of intricate hand interactions in real-world scenarios. As such tokenization indicates the relative relationship of two hands, it also supports more effective feature fusion. To this end, we further develop a Spatio-temporal Interaction Reasoning (SIR) module to fuse hand tokens in 4D with attention and decode them into 3D hand meshes and relative temporal movements. The efficacy of our approach is validated on several benchmark datasets. The results on in-the-wild videos and real-world scenarios demonstrate the superior performances of our approach for interactive hand reconstruction. More video results can be found on the project page: https://4dhands.github.io.
Figure 1: The proposed method, 4DHands, can robustly recover interactive hand meshes and their relative movement from monocular inputs.
11footnotetext: Corresponding authors.
1 Introduction
Hand mesh recovery from monocular inputs is a fundamental technology for numerous advanced applications like virtual/augmented reality, human-computer interaction, robotics, and embodied AI. Despite the remarkable progress achieved by existing methods [27, 29, 53] in this field, challenges remain in addressing generalization under in-the-wild scenarios, capturing robust hand motions with temporal stability over time, and recovering plausible interactions between two hands.
With the developments in transformer architectures, recent state-of-the-art approaches [39] to monocular human/hand mesh recovery have shifted toward investigating the usage of simple and high-capacity models along with large-scale in-the-wild data. Despite the improvement in 2D alignment, existing approaches still can not robustly handle close interaction and reconstruct hand meshes in the 4D space with stable temporal movement. Moreover, existing approaches typically handle two-hand and single-hand scenarios differently. As the two-hand status frequently switches between close interaction and separation in real-world scenes, the inconsistent processing procedure inevitably leads to clear jittering artifacts and temporal instability.
In this paper, we identify the three major limitations in previous solutions: i) the lack of a unified solution to handle hand image inputs in various situations, ii) neglecting the positional relationship between the hands within images, and iii) unstable estimation of the temporal translation of hands especially during the status shift between interaction and separation. To address these issues, we proposed 4DHands, a robust solution to recover interactive hand meshes and their relative movements in 4D from monocular inputs. The key insight of our solution is to unify the feature learning of two hands in various scenarios and fully leverage their positional relationship. To achieve this, 4DHands adopts a transformer-based architecture and includes novel tokenization and 4D feature fusion strategies.
Specifically, we first propose a Relation-aware Two-Hand Tokenization (RAT) method to embed positional relation information of each hand into its tokens. Such a tokenization method brings the following benefits threefold. i) RAT is not reliant on the co-occurrence of two hands and can thus handle both single-hand and two-hand inputs in a unified manner. ii) RAT takes the overlapped status of two hands into account and adjusts the tokens accordingly, which allows the network to learn a more concentrated distribution of the hand tokens regardless of the various scales and positions. iii) RAT explicitly encodes the relative positions of two hands into the tokens, facilitating the reconstruction of close and intricate interaction in real-world scenarios. After obtaining the image tokens, we further develop a Spatio-temporal Interaction Reasoning (SIR) module to infer hand poses and the relative positions between the two hands. Thanks to the proposed tokenization method RAT, SIR can fuse both the spatial and temporal information more effectively as the hand tokens indicate the relative relationship of two hands, thereby enhancing the robustness of hand mesh and temporal movement recovery in the 4D space.
In summary, the main contributions of this work can be listed as follows:
  • We introduce a Relation-aware Two-Hand Tokenization (RAT) in our transformer. By embedding positional relation information in the tokenization process, RAT enables a generic and informative process of hand inputs across diverse scenarios, regardless of whether the two hands are interacting or separated.
  • We propose a Spatio-temporal Interaction Reasoning (SIR) module to fuse 4D hand features. The collaboration of RAT and SIR further enhances the stability of hand mesh and temporal movement recovery in the 4D space.
  • Our final solution, 4DHands, achieves robust 4D hand mesh recovery in real-world scenarios, serving as a versatile solution to handle hand inputs in various forms, whether the images contain single or two hands, in single frames or sequences.
2 Related Work
2.1 Monocular Hand Mesh Recovery
Recent years have witnessed great progress in hand mesh recovery from monocular RGB input. Following the advances in monocular human mesh recovery [20, 24, 23, 56, 55, 45], some researchers [1, 39, 37, 58, 2] use neural networks to regress the driving parameters of hand templates such as MANO [41] to recover hand meshes, while others prefer to directly estimate the vertex coordinates of hand meshes through GCN [14, 25, 44, 6, 4, 5] or transformer [30, 29].
Recently, there have also been some efforts to simultaneously recover both two hands of humans. However, two hands mesh recovery is much more challenging than one hand, due to larger pose distribution and more severe occlusion. Furthermore, compared to the large-scale single-hand datasets [39] with rich capture scenarios and postures, existing two-hand datasets are much fewer, and most of them are captured in limited environments.
A straightforward method for two-hand mesh recovery is to locate and crop each hand separately from the input image, and then convert the task into single-hand scenarios by processing two cropped hand images individually. This strategy is wildly used in full-body pose estimation frameworks [57, 19, 8, 42, 12, 35, 59]. However, this approach is still based on single-hand recovery network, which usually performs poorly under close interaction situations. Moreover, the relative translation between two hands is ambiguous when the two hands are processed separately.
2.2 Hand Mesh Recovery under Interaction
To better address two-hands pose estimation and mesh recovery in close interaction scenarios, some researchers [27, 54, 53, 63, 21, 34, 32, 28] used one image containing both hands as input. Zhang et al. [54] utilized 2d joints projection attention and context-aware refinement module to improve the pose accuracy. Li et al. [27] utilize transformer module [9] to learn implicit attention between two hands. Zuo et al. [63] propose a data-based prior of two interacting hands. Yu et al. [53] constructed the attention relationship between the two hands based on 2D hand center maps. Although the above approaches have produced promising results, there are still some limitations. Because these methods use a single image containing both hands as input, when the two hands are far apart, the area proportion of hands in the input image will be too small, which can reduce the accuracy of mesh recovery. Furthermore, their network design inherently lacks the ability to recover single-hand mesh and only can operate under two-hand scenarios.
Therefore, we propose a new network that is compatible with both single-hand recovery and two-hand recovery. We take one or two cropped hand images as input, enabling utilize large-scale single-hand datasets to boost two-hand recovery results. We also tokenize the 2D relative distance and overlap between two cropped hand images by a powerful transformer [48, 9] module, resulting in accurate relative root translation estimation and interacting hands mesh recovery.
2.3 Temporal Hand Mesh Recovery
Several methods have been proposed for hand mesh recovery from temporal sequence data [31, 17, 52, 13, 22, 7]. Hasson et al. [17] leverage the optical flow of nearby frames to constrain the motion of hands and objects by photometric supervision. SeqHAND [52] incorporates the convolution-LSTM layer to capture the sequential relationship between consecutive hand poses. Deformer [13] proposes a novel dynamic fusion module that explicitly deforms nearby frames with clear hand visibility for robust hand pose estimation from temporal input. VIBE [22] and TCMR [7] utilize the recurrent neural network to produce temporally consistent motion.
However, these methods rely solely on sequence data. Our approach introduces a Spatio-Temporal Interaction Reasoning (SIR) module, enabling our method to not only recover accurate two-hand interacting meshes from single-frame input but also produce smooth and stable 4D hand recovery from occluded and motion-blurred temporal input.
3 Method
Figure 2: Overview of our 4DHands framework. 4DHands is a transformer-based network, which takes monocular images as input and estimates two-hand meshes with their relative positions.
In this Section, we present the technical details of 4DHands. As illustrated in Fig 2, 4DHands is a transformer-based network, which takes hand images as input and estimates the hand meshes, as well as their relative position. In the proposed transformer, the key components include Relation-aware Two-hand Tokenization (RAT) and Spatio-temporal Interaction Reasoning (SIR). As will be discussed shortly, the tokenization and reasoning strategies make 4DHands a generic solution to handle hand images in various forms, regardless of the images with single or two hands, in a single frame or sequences.
3.1 Problem Formulation
Given a single-frame or sequential image inputs $I$, our goal is to reconstruct the hand meshes and their relative position. We leverage the widely used parametric hand model MANO [41]. MANO defines a function $\mathbf{M}(\theta,\beta)$ that takes two parameters pose $\theta\in\mathbb{R}^{48}$ and shape $\beta\in\mathbb{R}^{10}$ as input, returns 3D hand mesh $V\in\mathbb{R}^{778\times 3}$ and 3D joint keypoints $V\in\mathbb{R}^{21\times 3}$. The outputs of our model are MANO parameters of two hands: $\theta_{R},\theta_{L},\beta_{R},\beta_{L}$, by which we can regress meshes and joints of each hand, and a vector $\upsilon$ to represent the joint root distance of two hands. We define $\upsilon$ to be the location of the left root joint location when we subtract the joints of both hands from the right hand’s 3D root joint location. Note that our method can robustly handle both single-hand and two-hand inputs. When there is only a single hand visible in the images, our method is simply degraded to the estimation of the mesh of one hand.
3.2 Relation-aware Two-Hand Tokenization
Existing approaches to monocular hand mesh recovery typically take the cropped hand images as input, but neglect the hand relationship in the images. Besides, these approaches handle two-hand and single-hand scenarios in different manners, resulting in jittering artifacts when the two hands operate in interaction and separately. To handle these issues in our transformer-based architecture, we propose a Relation-aware Two-Hand Tokenization (RAT) method to take the relation of two hands into consideration and handle inputs in various scenarios. Following recent state-of-the-art practices in 3D human/hand mesh recovery [51, 15, 39], we utilize ViT [9] as the visual encoder to tokenize image patches. In contrast to these methods, we handle both the single-hand and two-hand inputs in a unified manner and consider embedding positional relation information of each hand into its tokens during the tokenization process.
To achieve this, we first locate each hand from inputs and center it into the corresponding left and right hand images. This strategy allows our method to learn a more concentrated distribution of the hand tokens regardless of the various scales and relative positions. Moreover, by tokenizing the images of two hands individually, our method can maintain the stability of inputs across various interaction scenarios, eventually enhancing the robustness of our solution. To compensate for the lack of relation information in the separated tokens, we further consider enriching the hand tokens with relative positional embedding information so that they can be aware of the relative position with respect to each hand and the input images.
Specifically, for each input image, we follow ViT [9] to segment it into $H\times W$ image patches and subsequently feed them into the ViT backbone, resulting in a $H\times W\times C$ vision token map. To enrich each token map with relationship information, we further calculate relation maps $\{\mathcal{R}_{R2L},\mathcal{R}_{L2R}\}$ and embed them into the final image tokens, so that they can contain both the patch-wise visual knowledge and patch-to-patch relation information. Basically, the relation maps are the composition of the relative distance map and the overlapping map. Without losing generalization, we describe the right-hand relation map $\mathcal{R}_{R2L}$ in detail as follows. The left-hand relation map $\mathcal{R}_{L2R}$ can be calculated similarly.
Relative Distance Map
The relative distance map is calculated based on the hand position in the original images. Let $\mathcal{B}=[c_{x},c_{y},s_{x},s_{y}]$ be the bounding box of one hand, $c_{x},c_{y}$ the x-coordinate and y-coordinate of the center position, and $s_{x},s_{y}$ the box scale lengths on x-axis and y-axis. We obtain a position map $\mathcal{C}\in\mathbb{R}^{H\times W\times 2}$, representing positions of the center in each image patch. For the patch $p(i,k)$ at the $i$-th column and the $k$-th row, its value is
$\mathcal{C}(p)=[c_{x}+\frac{(2i-W)\cdot s_{x}}{2W},c_{y}+\frac{(2k-H)\cdot s_{y}}{2H}].$ (1)
In our method, for each right hand and left hand, we first obtain their position maps $\mathcal{C}_{R},\mathcal{C}_{L}$ and then calculate a relative distance map $\mathcal{D}\in\mathbb{R}^{H\times W\times 2}$. The relative distance token at $p(i,k)$ of the right hand is formulated as
$\mathcal{D}_{R2L}(p)=\psi(\tau\cdot\frac{\mathcal{C}_{R}(p)-\mathcal{C}_{L}(p)}{s})$ (2)
where $s=(s_{x},s_{y})$ is the scale of right-hand bounding box, $\psi$ is an activation function, $\tau$ is a hyper-parameter used to adjust the activation distance. In this way, the relative distance map is activated when two hands are close. In our experiments, we use Sigmoid as the activation function $\psi$.
Overlapping Map
To explicitly represent the overlapped relation of two hands, we further introduce the overlapping map. Denoting $Inside(p(i,k),\mathcal{B})$ as a function that indicates whether the center of patch $p$ is inside the bounding box $\mathcal{B}$, we formulate the the overlapping map $\mathcal{O}\in\mathbb{R}^{H\times W\times 1}$ as
$\mathcal{O}_{R2L}(p)=\begin{cases}1,&Inside(p,\mathcal{B}_{L})\\ -1,&\text{otherwise}\end{cases}$ (3)
Given the above relative distance map and overlapping map, the relation map is calculated by first concatenating them in the last dimension and using an MLP to match the dimension with the image feature dimension $C$:
$\mathcal{R}_{R2L}=MLP(\mathcal{D}_{R2L}\oplus\mathcal{D}_{L2R}\oplus\mathcal{O}_{R2L}),$ (4)
where $\oplus$ denotes the concatenation operation. Finally, the relation map acts as the relative positional embedding and enhances the original ViT tokens $F_{R}$ by concatenating them as the relation-aware tokens $F^{+}_{R}$.
3.3 Spatio-temporal Interaction Reasoning
After obtaining the image tokens, we further leverage a Spatio-temporal Interaction Reasoning (SIR) module to infer the hand poses and the relative position between two hands. Specifically, we flatten and concatenate the tokens of two hands to a sequence $M^{*}\in\mathbb{R}^{2HW\times C}$. The token sequence is further fed into attention modules to fuse both the spatial and temporal information.
Adaptive Spatial Fusion
To enhance the adaptability of our model, we first employ a self-attention transformer to proceed with the token sequence $M^{*}$. As the relation-aware tokens are embedded with both visual and relation features, the self-attention module can adjust its weight adaptively based on visual and relative positional associativity and mitigate the interdependence of bilateral hand information. For instance, when relation-aware tokens indicate that the two hands are far away from each other, the self-attention can mainly interact with the tokens of each hand individually. The outputs from the self-attention module are further fed into a transformer decoder, resulting in a global feature $G$ corresponding to each hand.
Temporal Fusion
When dealing with the sequential inputs, the global features of each frame are concatenated along the temporal dimension to form the temporal feature $\{G_{t}\}_{t=1}^{T}$, which will be further proceeded by a temporal decoder. By fusing the temporal information, our model can predict the hand poses of intermediate frames more robustly in handling disturbances caused by object occlusions and dynamic blurring. Note that our method is also compatible with single-frame inputs. In this case, we stack the global features repetitively to obtain the feature $\{G_{t}\}$ and feed it into the temporal decoder.
After the feature fusion of both spatial and temporal dimensions, we can obtain the integrated features $G^{*}_{R}$ and $G^{*}_{L}$ of the right and left hand, respectively. Finally, these features are proceeded by an MLP-based regression network to estimate the MANO parameters of two hands and their relative position.
3.4 Losses
As our method predicts two hands in a disentangled way, we introduce single-hand losses and inter-hand losses to improve the accuracy of single-hand reconstruction and two-hand interaction.
3.4.1 maxMSE Loss
Following [13], we use the maxMSE Loss to adjust the weights of each joint to maximize the MSE loss. Given joints locations $\mathcal{P^{*}},\mathcal{P}$, the maxMSE loss is formulated as
$maxMSE(\mathcal{P},\mathcal{P^{*}})=\frac{\sum_{i=1}^{N}||\mathcal{P}_{i}-\mathcal{P^{*}}_{i}||^{4}}{\sum_{i=1}^{N}||\mathcal{P}_{i}-\mathcal{P^{*}}_{i}||^{2}}.$ (5)
With maxMSE loss, the joints or vertices with large errors, typically the vertices of fingertips, will be assigned larger loss weights.
3.4.2 Single-Hand Losses
[…]
$L_{mano}=||\tilde{\theta}-\theta||^{2}+||\tilde{\beta}-\beta||^{2}$ (6)
With the pose and shape parameters, we can obtain 3-D vertices $V$ and 3-D joint keypoints $J$. With accurate 3-D ground truth $\tilde{V}$ and $\tilde{J}$, we can supervise the output 3-D keypoints with 3-D loss using maxMSE
$L_{3D}=maxMSE(\tilde{V},V)+maxMSE(\tilde{J},J)$ (7)
With ground truth 3-D joint keypoints and camera parameters, we can get 2-D keypoints $\tilde{j}=\tilde{\pi}(\tilde{J})$ on the original image. The predicted 2-D keypoints are calculated with predicted camera parameter $\pi$. The 2-D projection loss is formulated
$L_{2D}=maxMSE(\pi(J),\tilde{j})$ (8)
3.4.3 Inter-Hand Losses
We utilize Inter-Hand losses in order to optimize the finger touching details of two hands. To calculate the relative distance between right hand keypoints and left hand keypoints, we relocate all keypoints to set the palm joint of right hand on the original coordinate. We denote $\phi_{j}(i,k)\in\mathbb{R}^{3}$ as the distance vector from the $i-th$ right hand joint to the $k-th$ left hand joint, and $\phi_{v}(i,k)\in\mathbb{R}^{3}$ as the distance vector of vertices, defined in the same way.
To supervise the interacting gesture, for ground truth joints and predicted joints, we calculate the distance from each right-hand joint to each left-hand joint, respectively. We formulate the joint-relation loss as:
$L_{jrel}=\sum_{i=1}^{N_{j}}\sum_{k=1}^{N_{j}}||\tilde{\phi}_{j}(i,k)-\phi_{j}(i,k)||^{2}$ (9)
where $N_{j}=21$ represents the number of joint keypoints. For a finer reconstruction in contact of hands, we supervise the close vertices distances as:
$L_{close}=\sum_{i=1}^{N_{v}}\sum_{k=1}^{N_{v}}\begin{cases}||\tilde{\phi}_{v}(i,k)-\phi_{v}(i,k)||^{2},&\text{if }||\tilde{\phi}_{v}(i,k)||^{2}\leq\alpha\\ 0,&\text{if }||\tilde{\phi}_{v}(i,k)||^{2}>\alpha\end{cases}$ (10)
where $N_{v}=778$ is the number of vertices, and $\alpha$ is a hyper-parameter, representing the distance threshold. We set $\alpha=0.005$ in experiments.
4 Experiments
4.1 Implementation details
We implement our model based on PyTorch [38]. We use a pretrained Vision Transformer following a huge design (ViT-h) as the backbone to encode image features. The resolution of input image is $256\times 192$ and the size of the feature map is $H=16,W=12,C=1280$. We use transformer encoders and decoders in our feature fusion modules. We train our model using AdamW optimizer with a linear learning rate decay schedule from an initial learning rate $1\times 10^{-5}$. The length of an image sequence is set to 9 with a gap of 5 frames to sample images in a consecutive video sequence.
4.2 Datasets
We use a training dataset with a mix up of interacting hand datasets and single-hand datasets. The interacting hand datasets are used to optimize the hand relations in cases that two hands are close. The single-hand datasets are used to optimize the accuracy of single-hand mesh prediction in cases where two hands are far away.
Interhand2.6M [34]is the first publicly available dataset of sequences of interacting hands with accurate annotations, which is widely used for training and testing in two-hand reconstruction methods. InterHand2.6M in 30 fps is for video-related research and the 5 fps version is a subset of the 30 fps version which removes the redundancy for single image research. In our method, we make use of 30 fps and 5 fps interacting two-hand (IH) data with human and machine (H+M) annotations and filter invalid data following [27]. We ultimately utilize 1,614K images for training and 601K images for testing.
Re:InterHand [36] has images with realistic and diverse appearances along with accurate GT 3D interacting hands. It has 738K video-based images in 3rd-person viewpoints. We take the whole dataset for training.
DexYCB [3] contains sequences of single hands interacting with objects. It consists of 1000 sequences with 582K images. Although it’s not a two-hand dataset, we use it as part of the training set and testing set to enhance our model’s performance in scenarios with object occlusion and motion blur.
Static Single Hand Datasets. To enhance the performance in wild videos, we use single-hand datasets that provide 3D hand annotations for training. Following the previous work [39], we use the datasets with a combination of FreiHAND [62], HO3D [16], MTC [50], RHD [61], H2O3D [16], DEX YCB [3], COCO WholeBody [18], Halpe [11] and MPII NZSL [43].
In-the-wild Datasets. We use in-the-wild datasets HIC [47], ARCTIC [10] and RenderIH [26] to evaluate the performance on various scenarios. HIC provides diverse hand-hand interacting and object-hand interacting sequences and contains 3D GT meshes of both hands. ARCTIC is a multi-view sequential dataset of two-hand interacting with daily use objects, involving occlusions and detailed interactive actions. RenderIH provides realistically rendered hands with a variety of poses, textures, backgrounds, and illuminations, which can simulate various scenarios in reality. For all in-the-wild datasets, we only use them for testing to evaluate the performance on unseen scenarios.
4.3 Metrics
We adopt Mean Per Joint Position Error (MPJPE) and Mean Per Vertex Position Error (MPVPE) and Area Under Curve (AUC) scores to measure the accuracy of the pose and shape of the estimated meshes, and Mean Relative-Root Position Error (MRRPE) to measure the accuracy of relative positions of two hands. Following the prior method [27], we perform root joint alignment and skeleton scaling when evaluating MPJPE. In addition, we use acceleration error Accel_E to evaluate the temporal consistency, following [22].
4.4 Qualitative Results
Our qualitative comparisons are shown in Figure 3. We demonstrate results of in-the-wild datasets ARCTIC and RenderIH, which are unseen in the training phrase. We compare with IntagHand[27] and ACR [53] in camera perspective and an arbitrary perspective to show more details. We show more results on realistic complex cross-hand interacting samples in 4. The results demonstrate the stability of 4D-Hands in real-world scenarios.
Figure 3: Qualitative Comparison. We compare 4DHands with state-of-the-art methods on in-the-wild datasets ARCTIC[10] and RenderIH[26]. ARCTIC provides two-hand object interacting images, and RenderIH provides interactive hands under different lighting conditions. We have demonstrated the results from both the camera perspective and an arbitrary perspective. The results show that our method outperforms the state-of-the-art methods in complex real-world environments.
Figure 4: Qualitative Results. We show our model’s results on complex cross-hand interactions in realistic scenarios to demonstrate its performance. For a detailed perception, views of multiple perspectives are demonstrated.
4.5 Comparison with state-of-the-art methods
To demonstrate the versatility and stability of 4DHands across various scenarios, we conducted comparative analyses with state-of-the-art methods in multiple domains, including temporal two-hand evaluation, frame-based two-hand evaluation, temporal single hand evaluation and in-the-wild dataset evaluation.
Comparison with Temporal Two Hand Methods. In Table 1 we compare the proposed 4D-Hands with previous temporal methods on Interhand2.6m 30fps. The results show that our method far outperforms the existing two-handed temporal models, with 45.20%(7.37 mm $vs.$ 13.45 mm) improvement in MPJPE and 24.05% (2.81 mm/$s^{2}$ $vs.$ 3.70 mm/$s^{2}$) improvement in Accel_E.
Comparison with Frame-based Two Hand Methods. To compare with state-of-the-art frame-based methods and demonstrate the performance in single frame input scenarios, we conduct experiments at Interhand2.6m 5fps by expanding single image to sequences. As can be observed at Table 2, even though we drop the temporal knowledge, our method still keep high performance and outperforms previous methods that were specially designed for single frame input scenarios, achieving 7.49 mm, 7.72 mm and 24.58 mm in MPJPE, MPVPE and MRRPE.
Table 1: Quantitative comparisons with frame-based and temporal methods on Interhand2.6m 30fps dataset.
[…]
SeqHand[52] 6.69 18.20 19.46
Zhao et.al[60] 3.70 13.45 13.91
Ours 2.81 7.37 7.58
Table 2: Performance comparisons on the Interhand2.6m 5fps dataset. We compare with the most recent methods.
[…]
InterWild[33] 11.12 13.01 29.29
IntagHand[27] 8.79 9.03 -
MEMAHand[49] 8.65 8.89 -
ACR[53] 8.41 8.53 -
Zuo et.al[63] 8.34 8.51 -
Ren et.al[40] 7.51 7.72 28.98
Ours 7.49 7.72 24.58
Comparison with Temporal Single Hand Methods. To further exploit the universality of our method, we conduct experiments on sequential single hand dataset Dex-YCB. We manage our input by flipping an input image as the opposite hand, and using far away bounding boxes to indicate the hands are not interacting. We report the results of Root-Aligned and Procrustes-aligned metrics. It can be observed in Table 3 that our method outperforms all existing single hand temporal methods in all metrics. It is noteworthy that our lead in Root-Aligned metrics is greater than those in Procrustes-aligned metrics (+38.63% $vs.$ +9.00% in MPJPE, +12.70% $vs.$ +0.89% in JAUC), indicating that we can predict accurate shapes and rotation angles without the assistance of ground truth alignment.
Table 3: Quantitative comparisons with other single hand temoral reconstruction methods on the Dex-YCB dataset.
[…]
VIBE[22] 16.95 67.5 6.43 87.1
TCMR[7] 16.03 70.1 6.28 87.5
Tu et al. [46] 19.67 62.5 7.27 85.5
MeshGraphormer[30] 16.21 69.1 6.41 87.2
Deformer[13] 13.64 74.0 5.22 89.6
Ours 8.37 83.4 4.75 90.4
Comparison on In-the-wild Dataset. We report the results on HIC dataset, which contains in-the-wild cross-hand data, to compare the performance at in-the-wild scenarios. We only use samples in which both hands are visible. In 4 we compare the results with IntagHand and InterWild, a SOTA method specifically designed for two-hand recovery in the wild. Comparing to the MPJPE results on Interhand2.6m, which is employed during training, the performance of IntagHand and InterWild deteriorates by 131.8% and 40.5%, respectively, while our method exhibits a more modest reduction of 24.43%. This demonstrates that our method has greater stability on unseen data.
Table 4: Performance comparisons on the HIC dataset. We compare with Intaghand [27] and InterWild [33].
[…]
Intaghand[27] 20.38 21.56 73.04
InterWild [33] 15.62 15.17 26.43
Ours 9.32 9.93 25.26
4.6 Ablation Study
ViT Baseline. We trained a baseline using ViT and a transformer decoder on full training set, cross-hand feature interaction is not employed on the baseline. We show the results in 5. Although ViT has high performance in detecting single hand, the occlusion caused by hands interaction reduces its predictive ability. Besides, without cross-hand knowledge, it separately predict positions of two hand, causing huge errors in relative position.
Effectiveness of Cross-hand Information. We experimented in adding cross-hand information in two ways: (1) Input holistic image including two hands(+Cross H). (2) Input separate images, each image contains a single hand(+Cross S). In both way we add attention transformer module to integrate cross-hand information, the results show that the last way achieves better accuracy. The main reason is that by separating two hands as input, the visual encoder focuses on extracting single hand features, allowing the ViT to learn more hand details when learning on large scale datasets.
Effectiveness of Relation-aware Tokenization. By removing RAT(+Cross S w/o RAT), the model lost global position information during inference, which means that the model could only infer two-hand relative position from visual information. In InterHand2.6m, the two hands are relatively close and the 2D scales of each hand are similar. Therefore, even without a global position information, the approximate positions of both hands can be inferred from the input image. We test cases when two hands are relatively far away and cases when 2D scales of each hand are different. As shown in 5, without RAT, the model performs poor in predicting two-hand relative position.
Effectiveness of Temporal Fusion. By adding temporal information interacting(+Cross S + Seq), our method can make sequential consistent prediction therefore infer reasonable poses encountering occlusions, and achieves improvement of 0.32 mm in MPJPE. In Figure 6 we visualize the predicted mesh sequences in the case of severe occlusion. It shows that the left hand is entirely occluded, the model without temporal fusion makes erroneous prediction. Our full model predicts a consistent sequence with assistant information from the past and the future.
Table 5: Ablation studies of the RAT and SIR modules on InterHand2.6M.
[…]
Baseline 10.94 11.35 137.53
+Cross H 8.78 9.01 28.65
+Cross S 7.69 7.92 24.79
+Cross S w/o RAT 7.81 8.00 25.81
+Cross S + Seq (Ours) 7.37 7.58 24.48
Figure 5: Ablation study of RAT. Two views of predicted hand meshes are visualized to demonstrate details of relative position. ’Full Model’ is our model with all modules. ’w/o RAT’ means that the position embedding is removed.
Figure 6: Ablation study on occlusion case. ’Full Model’ is our model with all modules. ’w/o Temporal Fusion’ means that the temporal integrating module is removed.
5 Conclusions
In this paper, we presented 4DHands, a robust solution for recovering interactive hand meshes and their relative movements from monocular inputs. 4DHands is a transformer-based method with novel tokenization and feature fusion strategies. To this end, we introduced the Relation-aware Two-Hand Tokenization (RAT) method, which embeds positional relation information into hand tokens. This allows our network to handle both single-hand and two-hand inputs, leveraging their relative positions to reconstruct detailed hand interactions in real-world scenarios. Moreover, this tokenization approach enhances feature fusion by indicating the relative positions of the hands. To further this goal, we developed a Spatio-temporal Interaction Reasoning (SIR) module that fuses hand tokens in 4D with attention, decoding them into 3D hand meshes and their temporal movements. The effectiveness of 4DHands is demonstrated through extensive evaluations on benchmark datasets, with results from in-the-wild videos and real-world scenarios showcasing its superior performance in interactive hand reconstruction.